# Network-Optimised Spiking Neural Network for Event-Driven Networking

### v3

@@ Discussion
Several limitations follow from our modeling choices. The arrival model uses compound Poisson shot noise with exponential smoothing, which captures burstiness at the timescales of typical telemetry but does not represent long-range dependence or heavy-tailed flow mixing in full generality. The experiments focus on single-queue observables and fixed topologies, and the queueing baselines largely reflect M/M/1-style behaviour. Joint adaptation with transport protocols, multi-queue schedulers and cross-layer policies remains to be explored. Training uses surrogate gradients and truncated BPTT windows aligned with the recovery time. Although the homotopy schedule and clipping rules stabilise optimisation, a more systematic study of learnability under aggressive thresholds and high coupling would be valuable.
Future work can proceed along three directions. First, richer traffic and topology models, including multi-tenant datacentre fabrics and radio access networks with mobility, would test the robustness of the stability proxy and the forecasting claims. Second, coupling NOS units with GNN or RL controllers in a principled hierarchical design would realise the complementary roles discussed in the introduction, with NOS providing fast, local reflexes and deep models orchestrating global policy. Third, hardware-in-the-loop studies on contemporary neuromorphic platforms would allow a direct comparison of energy and tail performance against software-only baselines. Together, these steps would clarify whether NOS can serve as a practical building block for AI-native networks in production rather than only as an analytic and experimental tool.
@@ Appendix A Spiking Neural Networks for Networking: Scope and Complementarity
Anomaly detection (e.g., DDoS, worm outbreaks) Event-driven detection of sudden bursts; low-latency alarms [Maciag2021, Baessler2022] Flow-level detection from large corpora; supervised and deep models [Ahmed2016, Jiang2023]
Routing optimisation Local adaptation in ad hoc or sensor settings using local rules; energy- and event-efficient deployment on neuromorphic hardware [Davies2018, Furber2014] Global path optimisation and traffic engineering with topology-wide objectives; GNN/RL controllers [XuTealSIGCOMM2023]
QoS prediction (throughput, delay) Edge-side reaction to micro-bursts under power limits (event-driven inference) [Davies2018] Predictive modelling from aggregates; offline/online regression and forecasting [Giannakou2020, Krishna2024, xu2023]
Traffic classification Temporal signatures (periodicity, jitter) via spike timing and event patterns [rasteh2022, plat2017] Feature-rich flow classification with labels using CNN/RNN/GNN [Nguyen2008, Zhao2021, basit2022]
Wireless resource allocation Local duty-cycling, sensing, and lightweight on-device decision-making under tight energy budgets [Furber2014] Centralised spectrum allocation and large-scale optimisation with RL/GNN [LuongDRL2019, Zhi2022, IQBAL2023]
Network resilience and fault detection Distributed detection of local failures and sudden state changes in streams [Maciag2021, Baessler2022] Correlating failures across topology; resilience modelling and recovery for SDN/CPS [Menaceur2023, SegoviaFerreiraACMCSUR2024, Liu2024]
@@
Edge devices and in-network elements often face tight power and memory budgets. Neuromorphic platforms (e.g., Loihi and SpiNNaker) report energy advantages for spiking workloads, which supports in situ decisions at low latency [Davies2018, SpiNNaker]. In contrast, data-centre controllers and core routers, with ample power and global context, typically profit from conventional ML and GNNs.
@@ B.1 Random Neural Network (Gelenbe)
The Random Neural Network (RNN) treats each neuron as an integer-valued counter driven by excitatory and inhibitory Poisson streams and yields a product-form steady state [gelenbe1989random]. Let $q_{i}=\Pr\{k_{i}>0\}$ denote the activity of neuron $i$. In steady state
@@
The queueing affinity and fixed points make calibration attractive. However, backbone and datacentre traces are overdispersed and often long-memory, so a stationary product form understates burst clustering and tail risk. This follows from the Poisson and exponential assumptions of the model, which wash out temporal correlation and refractoriness. Product-form results emphasise stationarity, whereas operations rely on transient detection and control. Parameters $r_{i}$ and $p_{ji}^{\pm}$ do not tie directly to per-link capacity, delay, or gating. Under strong excitation, the signal-flow equations can fail to admit a valid solution with $q_{i}<1$, leading to instability or saturation; when a solution exists it is unique [gelenbe1989random, GelenbeNECO1990].
@@ B.2 Hodgkin–Huxley
The Hodgkin–Huxley (HH) equations provide biophysical fidelity through voltage-gated conductances and reproduce spikes, refractoriness and ionic mechanisms [Hausser2000]. The membrane and current relations are
@@
$I_{\mathrm{Na}}$ $=\bar{g}_{\mathrm{Na}}\,m^{3}h\,(V-E_{\mathrm{Na}}),\qquad I_{\mathrm{K}}=\bar{g}_{\mathrm{K}}\,n^{4}\,(V-E_{\mathrm{K}}),\qquad I_{L}=\bar{g}_{L}(V-E_{L}),$ (11)
@@
Hodgkin–Huxley offers mechanistic fidelity and a broad repertoire of excitable behaviour, and recent results show that HH neurons can be trained end to end with surrogate gradients, achieving competitive accuracy with very sparse spiking on standard neuromorphic benchmarks [HHExtremeSparsity]. This confirms feasibility for learning and suggests potential efficiency from activity sparsity. For network packet-level control, the obstacles are practical rather than conceptual. The model comprises four coupled nonlinear differential equations with voltage-dependent rate laws, which typically require small timesteps and careful numerical treatment; in practice this raises computational cost and stiffness at scale, and training can fail without tailored integrators and schedules [HHExtremeSparsity]. There is also no direct mapping from biophysical parameters to queue occupancy, service rate, or link delay, so calibration from network telemetry lacks interpretability. Taken together, HH remains a benchmark for cellular electrophysiology [Hausser2000], but it is a poor fit for networking tasks that requires lightweight, semantically mapped states and trainable dynamics.
@@ B.3 Izhikevich
The Izhikevich model balances speed and diversity of firing patterns [Izhikevich2003]. Its dynamics and reset are
$\frac{dv}{dt}$ $=0.04v^{2}+5v+140-u+I,\qquad\frac{du}{dt}=a(bv-u),$ (13)
$\text{if }v\geq 30\,\mathrm{mV}:\quad$ $v\leftarrow c,\quad u\leftarrow u+d.$ (14)
@@ Appendix C Admissible parameter region and calibration conventions
Table 4: Admissible parameter region used in the experiments. Entries enforce $v\in[0,1]$, a bin width of 55 ms, and keep the stability condition from Lemma 2 satisfied with margin in operating regimes. Values were sampled within these regions; defaults are given elsewhere for reproducibility.
@@
$\alpha$ excitability scale in $f_{\mathrm{sat}}$ [0.4, 1.0][0.4,\,1.0] (dimensionless)
$\kappa$ saturation knee of $f_{\mathrm{sat}}$ [0.0,0.5, 2.0][0.0,0.5,\,2.0] (dimensionless)
$\beta$ linear excitability gain [0.10, 0.8][-0.10,\,0.8] (dimensionless)
$\gamma$ constant drive (baseline load) [0.00, 0.15][0.00,\,0.15] (in $v$ units)
$\lambda$ service/leak on $v$ [0.10, 0.30][0.10,\,0.30] (per bin [^2] Per-bin rates correspond to a $5$ ms bin width. Multiply by $200$ to convert to s-1.)
@@
$v_{\mathrm{th}}$ firing threshold [0.50, 0.68][0.50,\,0.68] (in $v$ units)
@@
$c$ post-event baseline level [0.0, 0.2][0.0,\,0.2] (in $v$ units)
$d$ recovery jump on event [0.1, 0.4][0.1,\,0.4] (in $u$ units)
$I_{0}$ drive offset (arrivals $\to\!I$) [0.08, 0.16][0.08,\,0.16] (in $v$ units)
@@
$\tau_{s}$ optional drive smoothing [0, 15][0,\,15] ms
$g$ coupling on $W$ (network) choose so $k_{\text{net}}=g\,\rho(W)\in[0,\,1.8]$
Shot noise arrival model for synthetic tests rate $\nu_{i}\in[10,\,50]$ Hz, amplitude $A\in[0.3,\,0.9]$
Delays link propagation $\tau_{ij}\in[0,\,25]$ ms
@@ Appendix D Design principles and engineering guidance
$\kappa_{\sigma}(t)\;=\;\kappa_{\sigma,0}+\bigl(\kappa_{\sigma,\mathrm{final}}-\kappa_{\sigma,0}\bigr)\,\min\!\left\{1,\frac{t}{T_{\mathrm{h}}}\right\},\;\;\kappa_{\sigma,0}\approx 1,\;\;\kappa_{\sigma,\mathrm{final}}\in[20,100].$ (17)
@@
$T_{\mathrm{BPTT}}\;\gtrsim\;\frac{c_{\!*}}{a+\mu},\qquad c_{\!*}\in[3,5]\ \text{(capturing about 95\%–99\%)},$ (18)
@@ D.1 Neuromorphic implementation details
$v_{q}$ $=\frac{1}{2^{n}}\,\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{v}{V}\,2^{n}\right),\,0,\,2^{m}-1\right),\qquad u_{q}\;=\;\frac{1}{2^{n}}\,\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{u}{V}\,2^{n}\right),\,0,\,2^{m}-1\right).$ (19)
@@
$v_{t+1}$ $=(1-S_{i}(t))\,\tilde{v}_{t+1}+S_{i}(t)\Big[c+\big(\tilde{v}_{t+1}-c\big)\,e^{-r_{\text{reset}}\Delta t}\Big],$ (20)
$u_{t+1}$ $=u_{t}+d\,S_{i}(t),$ (21)
@@
$\Delta_{\mathrm{net}}$ $>g\,\rho(W)\,f^{\prime}_{\mathrm{sat}}(v^{*})\bigl(\delta_{f^{\prime}}+\delta_{W}+\delta_{g}\bigr)\;+\;\varepsilon_{\mathrm{LUT}}\,\rho\,\mathbb{E}\{|v-c|\},$ (25)
@@ Appendix E Spike generation and resets
$-r_{\text{reset}}\,\sigma_{\kappa_{\sigma}}\!\bigl(v_{i}-v_{\mathrm{th}}(t)\bigr)\,\bigl(v_{i}-c\bigr),\;\;\sigma_{\kappa_{\sigma}}(x)=\frac{1}{1+e^{-\kappa_{\sigma}x}},\;\;r_{\text{reset}}>0.$ (28)
@@ Appendix G Noise sensitivity under stochastic drive
$\eta_{i}(t)$ $=\sum_{n=1}^{N_{i}(t)}A_{i,n}\,e^{-(t-t_{i,n})/\tau_{s}}\,H(t-t_{i,n}),\qquad N_{i}(t)\sim\mathrm{Poisson}(\nu_{i}t),$ (35)
@@
$\mathbb{E}[\eta_{i}]$ $=\nu_{i}\,\mathbb{E}[A_{i}]\,\tau_{s},\qquad\mathrm{Var}[\eta_{i}]=\nu_{i}\,\mathbb{E}[A_{i}^{2}]\,\frac{\tau_{s}}{2},\qquad\mathrm{Cov}_{\eta_{i}}(\tau)=\nu_{i}\,\mathbb{E}[A_{i}^{2}]\,\frac{\tau_{s}}{2}\,e^{-|\tau|/\tau_{s}}.$ (36)
@@
$H_{v}(0)$ $=\frac{a+\mu}{ab-(a+\mu)\bar{d}},\qquad\sigma_{v}^{2}\;=\;\frac{1}{2\pi}\int_{-\infty}^{\infty}\bigl|H_{v}(i\omega)\bigr|^{2}\,S_{\eta_{i}}(\omega)\,d\omega.$ (39)

### v4

@@ 5 Discussion
A few limitations follow from our modelling choices. The arrival model uses compound Poisson shot noise with exponential smoothing, which captures burstiness at typical telemetry timescales but does not represent long-range dependence or heavy-tailed flow mixing in full generality. The experiments focus on single-queue observables and fixed topologies, and the queueing baselines largely reflect M/M/1-style behaviour. Joint adaptation with transport protocols, multi-queue schedulers, and cross-layer policies remains open.
6 Conclusion
We introduced Network-Optimised Spiking (NOS), a two-state event-driven unit designed for delay-coupled, graph-structured telemetry with explicit queue semantics. NOS combines bounded excitability, explicit service and damping leaks, delayed graph-local coupling with optional per-link gates, and differentiable resets that support surrogate-gradient training. We provided node- and network-level stability conditions, including a scalar spectral proxy based on $g\,\rho(W)$ that separates topology from local dynamics, and we connected stochastic drive under telemetry smoothing to increased variability near stability boundaries. Empirically, under a strict label-free residual protocol, NOS improves early-warning quality and detection latency over compact learned baselines across canonical topologies, while retaining calibration handles that are interpretable to operators.
Looking ahead, three directions are immediate. Richer traffic and topology models, including multi-tenant datacentre fabrics and mobile radio access networks, would test robustness of the stability proxy and forecasting claims. Hierarchical designs that couple NOS units with GNN or RL controllers could combine fast local reflexes with global policy. Finally, hardware-in-the-loop studies on contemporary neuromorphic platforms can quantify energy, latency, and tail behaviour under realistic deployment constraints.
Impact Statement
NOS supports stability-aware, event-driven inference for delayed, graph-coupled streams, with parameters that can be calibrated from telemetry and checked against simple spectral margins. It can reduce detection latency and severe congestion events in resource-constrained deployments by computing mainly when activity changes.
@@ Appendix A Spiking Neural Networks for Networking: Scope and Complementarity
Anomaly detection (e.g., DDoS, worm outbreaks) Event-driven detection of sudden bursts; low-latency alarms [25, 4] Flow-level detection from large corpora; supervised and deep models [2, 18]
Routing optimisation Local adaptation in ad hoc or sensor settings using local rules; energy- and event-efficient deployment on neuromorphic hardware [7, 8] Global path optimisation and traffic engineering with topology-wide objectives; GNN/RL controllers [35]
QoS prediction (throughput, delay) Edge-side reaction to micro-bursts under power limits (event-driven inference) [7] Predictive modelling from aggregates; offline/online regression and forecasting [13, 3, 34]
Traffic classification Temporal signatures (periodicity, jitter) via spike timing and event patterns [30, 29] Feature-rich flow classification with labels using CNN/RNN/GNN [28, 40, 33]
Wireless resource allocation Local duty-cycling, sensing, and lightweight on-device decision-making under tight energy budgets [8] Centralised spectrum allocation and large-scale optimisation with RL/GNN [24, 41, 16]
Network resilience and fault detection Distributed detection of local failures and sudden state changes in streams [25, 4] Correlating failures across topology; resilience modelling and recovery for SDN/CPS [26, 31, 23]
@@
Edge devices and in-network elements often face tight power and memory budgets. Neuromorphic platforms (e.g., Loihi and SpiNNaker) report energy advantages for spiking workloads, which supports in situ decisions at low latency [7, 9]. In contrast, data-centre controllers and core routers, with ample power and global context, typically profit from conventional ML and GNNs.
@@ B.1 Random Neural Network (Gelenbe)
The Random Neural Network (RNN) treats each neuron as an integer-valued counter driven by excitatory and inhibitory Poisson streams and yields a product-form steady state [11]. Let $q_{i}=\Pr\{k_{i}>0\}$ denote the activity of neuron $i$. In steady state
@@
The queueing affinity and fixed points make calibration attractive. However, backbone and datacentre traces are overdispersed and often long-memory, so a stationary product form understates burst clustering and tail risk. This follows from the Poisson and exponential assumptions of the model, which wash out temporal correlation and refractoriness. Product-form results emphasise stationarity, whereas operations rely on transient detection and control. Parameters $r_{i}$ and $p_{ji}^{\pm}$ do not tie directly to per-link capacity, delay, or gating. Under strong excitation, the signal-flow equations can fail to admit a valid solution with $q_{i}<1$, leading to instability or saturation; when a solution exists it is unique [11, 12].
@@ B.2 Hodgkin–Huxley
The Hodgkin–Huxley (HH) equations provide biophysical fidelity through voltage-gated conductances and reproduce spikes, refractoriness and ionic mechanisms [15]. The membrane and current relations are
@@
$\begin{split}I_{\mathrm{Na}}=\bar{g}_{\mathrm{Na}}\,m^{3}h\,(V-E_{\mathrm{Na}}),I_{\mathrm{K}}=\bar{g}_{\mathrm{K}}\,n^{4}\,(V-E_{\mathrm{K}}),\\ I_{L}=\bar{g}_{L}(V-E_{L}),\end{split}$ (11)
@@
Hodgkin–Huxley offers mechanistic fidelity and a broad repertoire of excitable behaviour, and recent results show that HH neurons can be trained end to end with surrogate gradients, achieving competitive accuracy with very sparse spiking on standard neuromorphic benchmarks [19]. This confirms feasibility for learning and suggests potential efficiency from activity sparsity. For network packet-level control, the obstacles are practical rather than conceptual. The model comprises four coupled nonlinear differential equations with voltage-dependent rate laws, which typically require small timesteps and careful numerical treatment; in practice this raises computational cost and stiffness at scale, and training can fail without tailored integrators and schedules [19]. There is also no direct mapping from biophysical parameters to queue occupancy, service rate, or link delay, so calibration from network telemetry lacks interpretability. Taken together, HH remains a benchmark for cellular electrophysiology [15], but it is a poor fit for networking tasks that requires lightweight, semantically mapped states and trainable dynamics.
@@ B.3 Izhikevich
The Izhikevich model balances speed and diversity of firing patterns [17]. Its dynamics and reset are
$\begin{split}\frac{dv}{dt}=0.04v^{2}+5v+140-u+I,\\ \frac{du}{dt}=a(bv-u),\end{split}$ (13)
$\begin{split}\text{if }v\geq 30\,\mathrm{mV}:\quad v\leftarrow c,\quad u\leftarrow u+d.\end{split}$ (14)
@@ Appendix C Admissible parameter region and calibration conventions
Table 4: Admissible parameter region used in the experiments. State $v$ is dimensionless; time is measured in bins with $\Delta t=5$ ms. We use the recovery coupling $\xi$ in (2) to map the (dimensionless) recovery state $u$ into a rate term, and fix $\xi=1$ in all experiments (absorbed into the definition of $u$). Per-bin rates convert to physical rates as $r_{\mathrm{phys}}=r/\Delta t$ (so 11 per bin =200=200 s-1). Values were sampled within these regions; defaults are given elsewhere for reproducibility.
@@
$\alpha$ excitability scale in $f_{\mathrm{sat}}$ [0.4, 1.0][0.4,\,1.0] (per bin)
$\kappa$ saturation knee of $f_{\mathrm{sat}}$ ($\kappa>0$) tested $\{0.5,\,2.0\}$ (dimensionless)
$\beta$ linear excitability gain [0.10, 0.8][-0.10,\,0.8] (per bin)
$\gamma$ constant drive (baseline load) [0.00, 0.15][0.00,\,0.15] (per bin)
$\lambda$ service/leak on $v$ [0.10, 0.30][0.10,\,0.30] (per bin)
@@
$\xi$ recovery-to-$\dot{v}$ coupling (maps $u\mapsto$ rate) fixed to $1$ (per bin)
@@
$v_{\mathrm{th}}$ firing threshold [0.50, 0.68][0.50,\,0.68] (dimensionless $v$ units)
@@
$c$ post-event baseline level [0.0, 0.2][0.0,\,0.2] (dimensionless $v$ units)
$d$ recovery jump on event (adds to $u$) [0.1, 0.4][0.1,\,0.4] (dimensionless $u$ units)
$I_{0}$ drive offset (arrivals $\to I$) [0.08, 0.16][0.08,\,0.16] (per bin)
@@
$\tau_{s}$ optional drive smoothing [0, 3][0,\,3] bins (equiv. [0, 15][0,\,15] ms)
$g$ coupling on $W$ (network) choose so $k_{\text{net}}=g\,\rho(W)\in[0,\,1.8]$ (dimensionless)
Shot noise arrival model for synthetic tests rate $\nu_{i}\in[10,\,50]$ Hz (mapped using $\Delta t=5$ ms), amplitude $A\in[0.3,\,0.9]$ (per bin)
Delays link propagation $\tau_{ij}\in[0,\,5]$ bins (equiv. [0, 25][0,\,25] ms)
@@
11footnotetext: Per-bin rates correspond to a 55 ms bin width. Multiply by $200$ to convert to s-1.
@@ Appendix D Design principles and engineering guidance
$\begin{split}\kappa_{\sigma}(t)\;=\;\kappa_{\sigma,0}+\bigl(\kappa_{\sigma,\mathrm{final}}-\kappa_{\sigma,0}\bigr)\,\min\!\left\{1,\frac{t}{T_{\mathrm{h}}}\right\},\\ \kappa_{\sigma,0}\approx 1,\;\;\kappa_{\sigma,\mathrm{final}}\in[20,100].\end{split}$ (17)
@@
$\begin{split}T_{\mathrm{BPTT}}\;\gtrsim\;\frac{c_{\!*}}{a+\mu},\\ c_{\!*}\in[3,5]\ \text{(capturing about 95\%–99\%)},\end{split}$ (18)
@@ D.1 Neuromorphic implementation details
$\begin{split}v_{q}=\frac{1}{2^{n}}\,\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{v}{V}\,2^{n}\right),\,0,\,2^{m}-1\right),\\ u_{q}\;=\;\frac{1}{2^{n}}\,\mathrm{clip}\!\left(\mathrm{round}\!\left(\frac{u}{V}\,2^{n}\right),\,0,\,2^{m}-1\right).\end{split}$ (19)
@@
$\begin{split}v_{t+1}=(1-S_{i}(t))\,\tilde{v}_{t+1}+S_{i}(t)\Big[c+\big(\tilde{v}_{t+1}-c\big)\,e^{-r_{\text{reset}}\Delta t}\Big],\\ u_{t+1}=u_{t}+d\,S_{i}(t).\end{split}$ (20)
@@
$\begin{split}\Delta_{\mathrm{net}}&>\;f^{\prime}_{\mathrm{sat}}(v^{*})\,\delta_{f^{\prime}}\;+\;g\,\rho(W)\,\bigl(\delta_{W}+\delta_{g}\bigr)\\ &\quad+\;\varepsilon_{\mathrm{LUT}}\,\rho(W)\,\mathbb{E}\{|v-c|\},\end{split}$ (24)
@@ Appendix E Spike generation and resets
$\begin{split}-r_{\text{reset}}\,\sigma_{\kappa_{\sigma}}\!\bigl(v_{i}-v_{\mathrm{th}}(t)\bigr)\,\bigl(v_{i}-c\bigr),\\ \sigma_{\kappa_{\sigma}}(x)=\frac{1}{1+e^{-\kappa_{\sigma}x}},\;\;r_{\text{reset}}>0.\end{split}$ (27)
@@ Appendix G Noise sensitivity under stochastic drive
$\begin{split}\eta_{i}(t)=\sum_{n=1}^{N_{i}(t)}A_{i,n}\,e^{-(t-t_{i,n})/\tau_{s}}\,H(t-t_{i,n}),\\ N_{i}(t)\sim\mathrm{Poisson}(\nu_{i}t),\end{split}$ (34)
@@
$\begin{split}\mathbb{E}[\eta_{i}]=\nu_{i}\,\mathbb{E}[A_{i}]\,\tau_{s},\qquad\mathrm{Var}[\eta_{i}]=\nu_{i}\,\mathbb{E}[A_{i}^{2}]\,\frac{\tau_{s}}{2},\\ \mathrm{Cov}_{\eta_{i}}(\tau)=\nu_{i}\,\mathbb{E}[A_{i}^{2}]\,\frac{\tau_{s}}{2}\,e^{-|\tau|/\tau_{s}}.\end{split}$ (35)
@@
$\begin{split}H_{v}(0)=\frac{a+\mu}{ab-(a+\mu)\bar{d}},\\ \sigma_{v}^{2}\;=\;\frac{1}{2\pi}\int_{-\infty}^{\infty}\bigl|H_{v}(i\omega)\bigr|^{2}\,S_{\eta_{i}}(\omega)\,d\omega.\end{split}$ (38)